integrate the following with respect to x up to a constant C:

  • 5x4- 4x3 + 2x - 3
  • 3sin x - 2cos x
  • (x3 + x)cos 5x
  • e2x + sin3x
1
Expert's answer
2021-06-18T10:52:03-0400

"1) \int (5x^4 - 4x^3 +2x-3)\,dx =x^5 - x^4+x^2-3x +C \\
2) \int (3\sin(x)-2\cos(x))\,dx =-3\cos(x)-2\sin(x) +C\\
3) \int (x^3+x)\cos(5x)\,dx= \\ = (x^3+x)\frac{1}{5}\sin(5x)-\int\frac{1}{5}(3x^2+1)\sin(5x)\,dx =\\
= (x^3+x)\frac{1}{5}\sin(5x) + \frac{1}{25}(3x^2+1)\cos(5x)-\frac{1}{25}\int6x\cos(5x)\,dx=\\
=(x^3+x)\frac{1}{5}\sin(5x) + \frac{1}{25}(3x^2+1)\cos(5x)-\frac{1}{125}6x\sin(5x)+\\+\frac{6}{5}\int\sin(5x)\,dx=(x^3+x)\frac{1}{5}\sin(5x) + \frac{1}{25}(3x^2+1)\cos(5x)-\\
-\frac{1}{125}6x\sin(5x)-\frac{6}{625}\cos(5x)+C = \\
=(\frac{x^3}{5}+\frac{19x}{125})\sin(5x)+(\frac{3x^2}{25}+\frac{19}{625})\cos(5x)+C\\
4)\int(e^{2x}+\sin(3x))\,dx=\frac{1}{2}e^{2x}-\frac{1}{3}\cos(3x)+C"